8. The relationship between marginal and average costs Consider the following scenario to understand the relationship between marginal and average values. Suppose Eric is a professional basketbali player, and his game log for free throws can be summarized in the following table. Fill in the columns with Erics free-throw percentage for each game and his overall free-throw average after each game. Percentage 75 75 2/8EEEEE8/16 2/4 8/10 8/10 10/20 18/30 26/40 On the following graph, use the orange points (square symbol) to plot Erics free-throw percentage for each game individuaily, and use the green points (triangle symbol) to plot his overall average free-throw percentage atter each game. Note: Piot your Game Free-Throw Percentage

media%2F361%2F3613e76b-182f-419a-91b6-e5

media%2Fc42%2Fc42cb37b-7008-4ba0-96b3-37

I could really use some help.
Answer #1

Game

Game Result

Total

Game Free-Throw %

Average Free-Throw %

1

6/8

6/8

75

75

2

2/8

8/16

2/8 x 100 = 25

8/16 x 100 = 50

3

2/4

10/20

2/4 x 100 = 50

10/20 x 100 = 50

4

8/10

18/30

8/10 x 100 = 80

18/30 x 100 = 60

5

8/10

26/40

8/10 x 100 = 80

26/40 x 100 = 65

Game Free-Throw Percentage Average Free Throw Percentage

a) rising

b) falling

c) rising

d) at its minimum